Mathematics 244 – Linear Algebra Definitions Quiz 2 – February 2, 2007

Your Name:

A) (5) Define: The set S = {v_{1}, v_{2}, …, v_{k}} of vectors in R^{n} is linearly independent.

B) (5) True or false: If v is in Span({v_{1}, …, v_{k}}), then the set {v, v_{1}, …, v_{k}} must be linearly dependent. Explain (briefly!)